11 – 5x + 3x + 4x =18
11 – 5x + 3x + 4x =18
11 – 5x + 3x + 4x =18

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

11 – 5x + 3x + 4x = 18

11 – 5x + 7x = 18

2x = 18 -11

x = 7/2

Thanks